Miller computer co. will manufacture 28,000 laptops in a year during one shift. there are fixed manufacturing costs of $2,000,000 per year and labor costs $9,200,000. miller computer co. would like to be able to manufacture 56,000 laptops next year and is considering adding a second shift. the unit labor cost for the second shift would be 25% higher than the day shift, but the total fixed manufacturing costs would increase only to $2,400,000.

a) what is the manufacturing cost per unit of the single shift model?

b) what is the average manufacturing cost per unit if miller computer co. runs 2 shifts?
Answer #1

Answer a) Miller computer co. will manufacture = 28,000 laptops in a year during one shift.

Fixed manufacturing costs = $2,000,000 per year

labor costs = $9,200,000.

So, Total Cost = Fixed Cost + Variable Cost

= $2,000,000 + $9,200,000

= $11,200,000

So, manufacturing cost per unit of the single shift model = Total Cost / Total number of units produced in the single shift

= 11,200,000 / 28,000

= $400

b) For the second shift:

Total fixed manufacturing costs = $2,400,000

Unit labor cost for the second shift =25% higher than the day shift

= $9,200,000 + ($9,200,000 * 25/100)

= $9,200,000 + $2,300,000

= $11,500,000

Total Cost only for the second shift = $2,400,000 + $11,500,000

= 13,900,000

Manufacturing cost per unit of the second shift model = Total cost for the second shift / Number of units produced in the second shift

= 13,900,000 / (56,000 - 28,000)

= 13,900,000 / 28,000

= $496.43 (rounding off to two decimal places)

Hence, the average manufacturing cost per unit if miller computer co. runs 2 shifts = (manufacturing cost per unit of the single shift model + Manufacturing cost per unit of the second shift model) / 2

= ($400 + $496.43) / 2

= $896.43 / 2

= $ 448.22 (rounding off to two decimal places)